Pixel:

Your health insurance premium is $4211 annually. If your employer pays 80% and you are paid monthly, how much is deducted from each monthly check?

Pixel:

4211 x .8 = 3368.80 4211 - 3368.80 = 842.20 842.20 / 12 = 70.1833....

TheSmartOne:

Looks good A way to do it faster, like I mentioned earlier, is to instead just multiply by 0.2 instead It's faster than multiplying by 0.8 and then subtracting it 4211 * 0.2 = 842.20